A pendulum consists of a 0.5kg metal ball suspended on a 0.50m string. The ball is pulled horizontally and up a total vertical distance of 0.10m. It is then released. At the bottom of the arc, the mechanical energy of the ball was determined to be 0.491 J. What was the speed of the ball at the bottom of its arc.
Answer #1

we know that

total mechanical energy = PE + KE

at bottom of arc , the ball has no potential energy.

it has KE = (1/2)*m*v^2

so mechanical energy = KE

0.491 = (1/2)*0.5*v^2

v = 1.4 m/s